3.2If a R200 billion increases in investment spending creates R200 billion of new income in the first round of the multiplier process and R160 billion in the second round. Calculate:
a. the marginal propensity to consume (MPC).
b. the value of the expenditure multiplier in this closed economy.
3.3 Assuming a private closed economy whereby the marginal propensity to consume is 0.9 and investment spending decreases by R1000 billion. What will be the change on equilibrium GDP?
1
Expert's answer
2020-04-29T09:25:27-0400

a. The marginal propensity to consume (MPC).

The MPC will be computed as using the following formula:

"\dfrac{\Delta Y}{\Delta I}=\dfrac{360}{200}=\dfrac{1}{1-MPC}"

Therefore,

"\dfrac{1}{MPS}=1.8"


"MPS=0.56"

Therefore, the MPC will be:

"MPC+MPS=1"


"MPC=1-0.56"


"MPC=0.46"

b. the value of the expenditure multiplier in this closed economy. 

The expenditure multiplier is estimated as follows:

"\text{Expenditure multiplier}=\dfrac{1}{MPC}"

"\text{ Thus Expenditure multiplier}=\dfrac{1}{0.46}=2.17"


Assuming a private closed economy whereby the marginal propensity to consume is 0.9 and investment spending decreases by R1000 billion. What will be the change on equilibrium GDP?

"\dfrac{\Delta Y}{\Delta I}=\dfrac{1}{1-MPC}"


"\dfrac{\Delta Y}{-1000}=\dfrac{1}{1-0.9}"


"\dfrac{\Delta Y}{-1000}=10"


"\Delta Y=10*-1000=-10000"

The equilibrium GDP will decrease with 10,000 billion as a result of 1000billion decrease in national investment spending.